# Re-parameterized Low-rank Prompt: Generalize a Vision-Language Model within 0.5K Parameters

### v2

@@ 4.2.2 Full-rank Initialization
In the field of tuning vision-language models, existing works have confirmed that the initialization method of prompts is quite important. For example,  [55, 54] adopt a hand-crafted template as the initial point of the text prompts, and  [24] copies the parameters in the text or image class token to initialize the prompts of the corresponding branch. If we just do a random initialization, the overall performance of such existing methods would drop severely. See Sec. 5.2 for more details. In other words, it would be helpful if we could take advantage of a good initial point.
@@
To solve such a problem, we add a concurrent branch of full-rank prompts $P_{fr}\in\mathbb{R}^{n\times d}$ along with the proposed low-rank prompts as shown in Fig. 3. Naturally, $P_{fr}=P_{init}$. And we randomly sample $P_{A}$/$P_{B}$ from a Gaussian distribution in which $\mu=0$ and $\sigma\rightarrow 0$. A small $\sigma$ here could avoid constant initialization and enrich the update paths. Notably, $P_{fr}$ is kept frozen during the whole adaptation, and thus it would not increase the number of tunable or stored parameters.
@@ 4.2.3 Regularization
As discussed in Sec. 3, existing works have shown that proper regularization would significantly improve the generalization ability. Therefore, to alleviate overfitting and catastrophic forgetting, we put a Dropout layer with drop ratio $p$ after the low-rank branch as displayed in Fig. 3.
Therefore, the input prompt of the feature encoder is
$P=P_{fr}+\mbox{Dropout}(P_{lr},p)$ (11)
Finally, we have
$P=P_{init}+\mbox{Dropout}(P_{A}P_{B},p)$ (12)
@@ 4.3 Efficiency Analysis
The whole adaptation process of a pre-trained vision-language model can be divided into three parts: training, storage and inference. In this subsection, we will analyze the efficiency of CLP in each part separately.
@@
Inference Before inference, we first load $P_{init}$, $P_{A}$ and $P_{B}$ from disk to memory. Noticing that Dropout is exactly an identity layer in the inference mode, we could pre-calculate the equivalent $P$ by
$P=P_{init}+P_{A}P_{B}$ (13)
and just keep $P$ in the memory. For inference, we directly use $P$ as the input prompts, and thus the inference cost is the same as classic prompt tuning. Some existing methods add complex bridges between the isolated parameters to earn extra improvements, e.g. CoCoOp [55]. There ain’t no such thing as a free lunch. They would face slower speed and huge memory occupation in inference time.
4.4 Advantages of RLP
  • Efficiency and Effectiveness To our knowledge, we are the first one to reach comparable or even better performance with state-of-the-art methods using a fairly small number of parameters, i.e. 516. RLP reaches an excellent trade-off between efficiency and effectiveness.
  • Simplicity Replacing the classic prompt by RLP just needs to modify several lines of code. The design of RLP merely changes the structure of prompts. No special loss nor extra module is introduced.
  • Robustness RLP is relatively capable of anti-disturbance. As in Tab. 3, RLP could maximally reserve its knowledge from domain shift.
  • Plug and Play In a classic transformer model, RLP only affects its prompts, which are very basic and necessary components in modern transformer-based architectures. Such characteristic enables us to plug RLP into most of the existing prompt tuning methods by simply replacing several lines of code.
@@ 5 Experiments
Table 2: Comparisons with latest methods in base-to-new generalization. H: harmonic mean [45]. Overall, RLP shows the strongest performance compared with other lastest tuning methods. RLP reaches highest average harmonic mean accuracy and wins on 7 of 11 datasets with far more fewer parameters.
(a) Average over 11 datasets.
#params Base New H
CLIP 0K 69.34 74.22 71.70
CoOp 2.1K 82.69 63.22 71.66
CoCoOp 35.4K 80.47 71.69 75.83
Adapter 525.5K 82.62 70.97 76.35
ProGrad 8.2K 82.79 68.55 75.00
RLP 0.5K 79.86 74.17 76.91
(b) ImageNet.
Base New H
CLIP 72.43 68.14 70.22
CoOp 76.47 67.88 71.92
CoCoOp 75.98 70.43 73.10
Adapter 76.53 66.67 71.26
ProGrad 77.03 68.80 72.68
RLP 76.07 70.87 73.37
(c) Caltech101.
Base New H
CLIP 96.84 94.00 95.40
CoOp 98.00 89.81 93.73
CoCoOp 97.96 93.81 95.84
Adapter 98.20 93.20 95.63
ProGrad 98.50 91.90 95.09
RLP 97.90 95.03 96.45
(d) OxfordPets.
Base New H
CLIP 91.17 97.26 94.12
CoOp 93.67 95.29 94.47
CoCoOp 95.20 97.69 96.43
Adapter 94.40 94.10 94.25
ProGrad 94.40 95.10 94.75
RLP 95.17 97.77 96.45
(e) StanfordCars.
Base New H
CLIP 63.37 74.89 68.65
CoOp 78.12 60.40 68.13
CoCoOp 70.49 73.59 72.01
Adapter 77.13 69.23 72.97
ProGrad 79.00 67.93 73.05
RLP 70.00 75.10 72.46
(f) Flowers102.
Base New H
CLIP 72.08 77.80 74.83
CoOp 97.60 59.67 74.06
CoCoOp 94.87 71.75 81.71
Adapter 97.70 70.83 82.13
ProGrad 96.27 71.07 81.77
RLP 94.80 74.97 83.72
(g) Food101.
Base New H
CLIP 90.10 91.22 90.66
CoOp 88.33 82.26 85.19
CoCoOp 90.70 91.29 90.99
Adapter 90.40 90.40 90.40
ProGrad 90.17 89.53 89.85
RLP 90.73 91.83 91.28
(h) FGVCAircraft.
Base New H
CLIP 27.19 36.29 31.09
CoOp 40.44 22.30 28.75
CoCoOp 33.41 23.71 27.74
Adapter 39.57 32.27 35.55
ProGrad 42.63 26.97 33.04
RLP 35.17 35.57 35.37
(i) SUN397.
Base New H
CLIP 69.36 75.35 72.23
CoOp 80.60 65.89 72.51
CoCoOp 79.74 76.86 78.27
Adapter 81.67 73.93 77.61
ProGrad 80.70 71.03 75.56
RLP 79.47 77.70 78.57
(j) DTD.
Base New H
CLIP 53.24 59.90 56.37
CoOp 79.44 41.18 54.24
CoCoOp 77.01 56.00 64.85
Adapter 80.47 52.23 63.35
ProGrad 76.70 46.67 58.03
RLP 74.27 60.33 66.58
(k) EuroSAT.
Base New H
CLIP 56.48 64.05 60.03
CoOp 92.19 54.74 68.69
CoCoOp 87.49 60.04 71.21
Adapter 86.93 64.20 73.86
ProGrad 91.37 56.53 69.85
RLP 81.57 61.67 70.23
(l) UCF101.
Base New H
CLIP 70.53 77.50 73.85
CoOp 84.69 56.05 67.46
CoCoOp 82.33 73.45 77.64
Adapter 85.80 73.63 79.25
ProGrad 83.90 68.50 75.42
RLP 83.33 75.03 78.97
To verify the effectiveness of the proposed method, we evaluate our method and make comparisons with the latest state-of-the-art methods in terms of the following problems:
  1. 1.
    Base-to-new generalization, which models are trained with base classes and evaluated on both base and new classes.
  2. 2.
    Domain generalization, which aims to show generalization to the domain shift, especially for out-of-distribution data.
  3. 3.
    Cross-dataset transfer, which aims to see if the method has the potential to transfer beyond a single dataset. It is a much more challenging problem because the fundamentals can be totally changed across different datasets.
  4. 4.
    Few-shot learning, which aims to evaluate the adaptation performance of the model to extract knowledge from a dataset whose samples are very few, e.g. 1, 2, 4, 8 or 16 samples.
All methods, including the baseline methods, in our experiments, are initialized with the same CLIP weights provided by the open-source CLIP [37] [^1] https://github.com/openai/CLIP.. In order to give details about the experimental setup, we first introduce the used datasets, the compared baseline methods, and the implementation details.
Datasets
Following previous work [54, 55], we leverage 11 image recognition datasets to verify the effectiveness of the proposed method for both the base-to-new generalization task. These datasets include two datasets for the generic object classification, i.e., ImageNet [5] and Caltech101 [8], five datasets for the fine-grained classification, i.e., OxfordPets [35], StanfordCars [23], Flowers102 [34], Food101 [2] and FGVCAircraft [33], one dataset for the scene recognition, i.e., SUN397 [46], one dataset for the action recognition, i.e., UCF101 [43], one dataset for the texture classification, i.e., DTD [4], and one dataset for the satellite imagery recognition, i.e., EuroSAT [12].
Competitors
  1. 1.
    CLIP [37] (ICML 2021): CLIP is a strong baseline vision-language model that is pre-trained on a large number of image-text pairs from the web by learning a contrastive objective. CLIP enables strong zero-shot adaptation ability on various downstream tasks by using fixed text prompts, i.e. a photo of a.
  2. 2.
    CoOp [54] (IJCV 2022): CoOp replaces the fixed text prompts in CLIP with tunable text prompts to improve the adaptation ability of the vision-language model. CoOp shows excellent performance in few-shot situations.
  3. 3.
    CoCoOp [55] (CVPR 2022): CoCoOp replaces the isolated tunable text prompts in CoOp with conditional text prompts, which receive extra gradients from the image features besides text features. CoCoOp largely improves the generalization ability of the vision-language model, getting good results on base-to-new generalization and domain adaptation.
  4. 4.
    CLIP-Adapter [10] (arxiv): CLIP-Adapter adopts the thoughts of classic Adapter [15] to use serial linear layers and activation functions to adapt for downstream tasks. It is simple yet effective in few-shot learning.
  5. 5.
    ProGrad [56] (ICCV 2023): ProGrad only updates the text and image prompts whose gradient are aligned (or non-conflicting) to the general knowledge, which is represented as the optimization direction offered by the pre-defined prompt predictions. Such regularization helps it finish good adaptation and generalization.
Table 3: Comparisons with latest methods in domain generalization. RLP gets comparable or even better results with the latest state-of-the art methods with much fewer parameters, showing excellent robustness for domain shift.
Source Target
#params ImageNet
-V2
-Sketch
-Adversarial
-Rendition
Average
CLIP 0K 66.73
60.83
46.15
47.77
73.96
57.18
CoOp 2.1K 71.51
64.20
47.99
49.71
75.21
59.28
CoCoOp 35.4K 71.02
64.07
48.75
50.63
76.18
59.91
Adapter 525.5K 69.33
62.53
47.67
49.17
75.42
58.70
ProGrad 8.2K 72.24 64.73
47.61
49.39
74.58
59.07
RLP 0.5K 70.80
63.95
49.07 50.97 77.19 60.30
Table 4: Results in the cross-dataset transfer setting. RLP gives the highest accuracy on 6 of 10 datasets, and slightly outperforms CoCoOp by a small margin on average though the source accuracy of RLP is a bit lower. Such result well demonstrates that RLP could maximally extract general and data-agnostic knowledge from given images.
Source Target
@@
CoOp 2.1K 71.51 93.70 89.14 64.51 68.71 85.30 18.47 64.15 41.92 46.39 66.55 63.88
CoCoOp 35.4K 71.02 94.43 90.14 65.32 71.88 86.06 22.94 67.36 45.73 45.37 68.21 65.74
Adapter 525.5K 69.33 93.43 88.87 64.40 70.27 85.63 24.67 65.80 44.90 47.70 66.00 65.17
RLP 0.5K 70.57 94.20 90.50 67.17 71.27 86.07 23.83 67.60 46.73 42.10 68.93 65.84
Training Details
For RLP, we conduct a grid search to find the optimal hyper-parameters based on the configuration of CoOp. We chose CoOp, the most lightweight prompt-tuning method, as well as the weakest method as our foundational model. In Sec. 5.1, we set the rank $r=1$ in RLP for all the settings in Sec. 5.1. The total number of tunable parameters in RLP is $4*1+1*512=516$, which is merely around 0.25x of that in CoOp. In Sec. 5.2, we will show how to turn such a weak baseline CoOp into an effective and efficient method step by step, only with the support of such $516$ parameters.
5.1 Main Results
5.1.1 Base-to-new Generalization
Following previous works [55, 54], for each dataset, we split its classes equally into two non-overlapping groups, i.e., one as base classes and the other as new classes. We train all models on the base classes and perform a base/new evaluation on the base/new classes.
The results are shown in Tab. 2. Overall, RLP still shows the strongest performance, and the superiority mainly relies on the improvement of new classes. In other words, RLP largely improves the generalization ability of CLIP. In particular, compared with our direct baseline CoOp, RLP gets $10.95$% accuracy gain on the new classes and $2.83$% accuracy drop on the base classes. Such a result tells us that our low-rank prompt and lightweight regularization design reach our initial purpose to make the model more generalizable, instead of overfitting every upcoming dataset.
Notably, compared with the latest lightweight method ProGrad, even if RLP only uses $<$20x of its parameters, RLP still outperforms it by a clear margin, which clearly demonstrates the excellent effectiveness and efficiency of our proposed method.
Figure 4: Average results of few-shot learning over 11 datasets.
@@ 5.1.1 Base-to-new Generalization
Low-rank prompts Full-rank init Dropout #params base new H

### v3

@@ 4.1.2 Special Initialization
In the field of tuning vision-language models, existing works have confirmed that the initialization method of prompts is quite important. For example,  [6, 7] adopt a hand-crafted template as the initial point of the text prompts, and  [29] copies the parameters in the text or image class token to initialize the prompts of the corresponding branch. If we just do a random initialization, the overall performance of such existing methods would drop severely. See Section 5.2 for more details. In other words, it would be helpful if we could take advantage of a good initial point.
@@
To solve such a problem, we add a concurrent branch of full-rank prompts $P_{fr}\in\mathbb{R}^{n\times d}$ along with the proposed low-rank prompts as shown in Fig. 2(b). Naturally, $P_{fr}=P_{init}$. And we randomly sample $P_{A}$/$P_{B}$ from a Gaussian distribution in which $\mu=0$ and $\sigma\rightarrow 0$. A small $\sigma$ here could avoid constant initialization and enrich the update paths. Notably, $P_{fr}$ is kept frozen during the whole adaptation, and thus it would not increase the number of tunable or stored parameters.
((a)) Comparison between DIP Tuning and Prompt Tuning.
((b)) DIP enjoys special initialization.
Figure 2: To switch from classic prompt tuning and to DIP tuning, just replace the ordinary prompts with DIPs. DIP introduces two small parameter matrices in shape $[n,r]$ and $[r,d]$ separately, and uses their product as an equivalent prompt in shape $[n,d]$. For the prompts with special initialization, e.g. a hand-crafted template “a photo of a” for the text prompts, we introduce a concurrent full-rank prompt branch along with the proposed low-rank prompts. By turning off the gradient of the newly added branch, we start training from a promising initial point, and the total number of tunable parameters or stored parameters will not increase as well. The Dropout layer could effectively regularize the update of low-rank prompts and alleviate overfitting and catastrophic forgetting. Dropout is a lightweight non-parametric layer and turns out to be an Identity layer in inference, resulting in negligible cost.
@@ 4.1.3 Regularization
As discussed in Section 2, existing works have shown that proper regularization would significantly improve the generalization ability. Therefore, to alleviate overfitting and catastrophic forgetting, we put a Dropout layer with drop ratio $p$ after the low-rank branch as displayed in Fig. 2(b).
Therefore, the input prompt of the feature encoder is $P=P_{fr}+\mbox{Dropout}(P_{lr},p)$. Finally, we have $P=P_{init}+\mbox{Dropout}(P_{A}P_{B},p)$.
@@ 4.2 Efficiency Analysis
The whole adaptation process of a pre-trained vision-language model can be divided into three parts: training, storage and inference. In this subsection, we will analyze the efficiency of DIP in each part separately.
@@
Inference Before inference, we first load $P_{init}$, $P_{A}$ and $P_{B}$ from disk to memory. Noticing that Dropout is exactly an identity layer in the inference mode, we could pre-calculate the equivalent $P$ by $P=P_{init}+P_{A}P_{B}$ and just keep $P$ in the memory. For inference, we directly use $P$ as the input prompts, and thus the inference cost is the same as classic prompt tuning. Some existing methods add complex bridges between the isolated parameters to earn extra improvements, e.g. CoCoOp [6]. There ain’t no such thing as a free lunch. They would face slower speed and huge memory occupation in inference time.
@@ 5 Experiments
To verify the effectiveness of the proposed method, we evaluate our method and make comparisons with the latest state-of-the-art methods in terms of the following settings in a wide range: base-to-new generalization, domain generalization, cross-dataset transfer and few-shot learning.
For more experimental details, please refer to the Appendix.
Table 2: Comparisons with latest methods in base-to-new generalization. H: harmonic mean [55]. In the lightweight methods, DIP+CoOp outperforms all the other competitors on new and harmonic mean accuracy. While in the heavy methods, DIP+MaPLe surpasses MaPLe [28] by a clear margin, and is strongest on base, new and harmonic mean accuracy, even with around only $\frac{1}{5}$ parameters compared with MaPLe.
Method #params Base New H
Lightweight
CLIP [1] 0K 69.34 74.22 71.70
CoOp [7] 2.1K 82.69 63.22 71.66
CoCoOp [6] 35.4K 80.47 71.69 75.83
Adapter [56] 525.5K 82.62 70.97 76.35
LoRA [37] 129.0K 84.30 67.33 74.86
ProGrad [52] 8.2K 82.79 68.55 75.00
DIP+CoOp 0.5K 80.32 74.73 77.42
Heavy
MaPLe [28] 3548K 82.28 75.14 78.55
DIP+MaPLe 741.9K 83.17 75.43 79.11
5.1 Main Results
5.1.1 Base-to-new Generalization
The average results over 11 datasets are shown in Tab. 2. For complete results, please refer to the Appendix. Overall, DIP shows the strongest performance, and the superiority mainly relies on the improvement of new classes. In other words, DIP largely improves the generalization ability of CLIP. In particular, compared with our direct baseline CoOp, DIP gets $10.95$% accuracy gain on the new classes and $2.83$% accuracy drop on the base classes. By adding DIP to the prompts as well as the linear layers of MaPLe, we earn a excellent 0.56% performance gain while reducing its parameters to 0.2×0.2$\times$0.2 × of the original parameters. Notably, compared with the latest lightweight method ProGrad, even if DIP only uses $<$20x of its parameters, DIP still outperforms it by a clear margin, which clearly demonstrates the efficiency and effectiveness.
Table 3: Comparisons with latest methods in domain generalization after tuned on ImageNet. DIP+CoOp shows excellent robustness when dealing with domain shift.
#params IN-V2 IN-Sketch IN-Adversarial IN-Rendition Average
CLIP 0K 60.83 46.15 47.77 73.96 57.18
CoOp 2.1K 64.20 47.99 49.71 75.21 59.28
CoCoOp 35.4K 64.07 48.75 50.63 76.18 59.91
Adapter 525.5K 62.53 47.67 49.17 75.42 58.70
LoRA 129.0K 62.37 42.43 38.40 68.97 53.04
ProGrad 8.2K 64.73 47.61 49.39 74.58 59.07
DIP+CoOp 0.5K 63.95 49.07 50.97 77.19 60.30
5.1.2 Domain Generalization
Then we follow CoCoOp [6] to use ImageNet, ImageNet-A, ImageNet-R, ImageNet-v2, and ImageNet-S to run domain generalization experiments to verify the robustness of DIP (DIP = DIP + CoOp, and the same applies thereafter, unless otherwise specified.). Shown in Tab. 3, on target datasets, DIP leads to better average accuracy compared with the latest methods, largely outperforming CLIP, CoOp, CLIP-Adapter and ProGrad with much fewer parameters.
Table 4: Results in the cross-dataset transfer setting. DIP+CoOp gives the highest accuracy on 6 of 10 datasets, and slightly outperforms CoCoOp on average. Such result well demonstrates that DIP could maximally extract general and data-agnostic knowledge from given images.
@@ 5.1.2 Domain Generalization
CoOp 2.1K 93.70 89.14 64.51 68.71 85.30 18.47 64.15 41.92 46.39 66.55 63.88
CoCoOp 35.4K 94.43 90.14 65.32 71.88 86.06 22.94 67.36 45.73 45.37 68.21 65.74
Adapter 525.5K 93.43 88.87 64.40 70.27 85.63 24.67 65.80 44.90 47.70 66.00 65.17
DIP+CoOp 0.5K 94.20 90.50 67.17 71.27 86.07 23.83 67.60 46.73 42.10 68.93 65.84
5.1.3 Cross-dataset Transfer
Finally, we follow CoCoOp [6] to conduct cross-dataset transfer evaluation. Results are shown in Tab. 4. Concentrating too much on the current dataset will absolutely cause overfitting and catastrophic forgetting problems, and finally lead to a severe drop in the performance on those unseen datasets. In this setting, DIP wins on 6 of 10 datasets and its average accuracy is also slightly better than the best competitor CoCoOp. Such result well demonstrates that DIP could maximally extract general and data-agnostic knowledge from given images compared with other prompt-based methods. Considering the huge difference in the parameter numbers, we could summarize that DIP is still the better choice.
5.1.4 Few-shot Learning
In this paragraph, we will show the experiment results of DIP in the few-shot learning setting. This setting is originated from CoOp. Seen from Fig. 4, DIP consistently outperforms zero-shot CLIP, CoOp, and CLIP-Adapter across all the shot numbers. Such results demonstrate the superiority of DIP in adaptation ability when there are few samples in downstream tasks.
Overall, in base-to-new generalization, domain generalization, cross-dataset transfer and few-shot learning, DIP can be fruitfully integrated into existing methods and consistently reaches state-of-the-art performance while enjoying extremely high parameter efficiency.
Figure 3: Few-shot learning Results.
Figure 4: Top: Effect of training epochs. Bottom: Effect of dropout ratios.
5.2 Analysis
@@
Decomposition Full-rank init Dropout #params base new H